Vetro has a customer who requires glass cut to a lower specification of 4.96 cm and an upper specification of 5.04 cm. Vetro's process averages 5 cm, but does have some variability. specifically, the standard deviation = 0.01 cm. What is Vetro's capability index for this customer? (Round to two decimal places)
Answer #1

Upper specification limit (USL) = 5.04

Lower specification limit (LSL) = 4.96

Mean = 5

Standard deviation (SD) = 0.01

Cpl = (Mean - LSL) / (3 x SD) = (5-4.96)/(3x0.01) = 0.04/0.03 = 1.33

Cpu = (USL - Mean) / (3 x SD) = (5.04-5)/(3x0.01) = 0.04/0.03 = 1.33

Process capability index (Cpk) = Minimum of (Cpu, Cpl)

= Minimum of (1.33 , 1.33)

= 1.33